The magnitudes of the current density and the diameters for wires 1 and 2 are given in the table. The current directions are indicated by the arrows.

Find the current I3 in wire 3.

Express your answer in amperes to two significant figures. Call current out of the junction positive and current into the junction negative.

The magnitudes of the current density and the diam

Wire Current density
(A/mm2)
Diameter
(mm)
1 2.1 1.7
2 3.9 2.4
Answer #1

current in wire 1 = I1 = current density * Area

I1 = 2.1*(pi*(1.7/2)^2) = 4.77 A

current in wire 2 = I2 = 3.9*(pi*(2.4/2)^2) = 17.65 A

kirchhoff law at juction point

I1 = I2 + I3

I3 = I1 - I2 = 4.77 - 17.65 = - 12.88 A

The negative sign indicates that I3 is entering the juction